OpenStudy (callisto):

What should I do to integrate this? \[\int \frac{4}{3+cosx} dx\]

OpenStudy (freckles):

Step 1: Multiply top and bottom by conjugate of the bottom

OpenStudy (amistre64):

what techniques you got at your disposal?

OpenStudy (freckles):

It will look juicy to us with a bottom with one term :)

OpenStudy (callisto):

Got it. So, would it be similar if I change 1 on the top to 4 and 1 in the bottom to 3?

OpenStudy (lgbasallote):

lol freckles juicy :P

OpenStudy (callisto):

Perhaps I shall rewrite the question first.

OpenStudy (callisto):

Latex fails :S This one: \[\int \frac{4}{3+cosx} dx\]

OpenStudy (lgbasallote):

is that the question now?

OpenStudy (freckles):

So this is a new problem you are saying? And yes you do multiply the conjugate of bottom on both top and bottom

OpenStudy (lgbasallote):

but pull out the constant first so it's not confusing :P

OpenStudy (lgbasallote):

i see this wont be as juicy as before...

OpenStudy (callisto):

\[4\int \frac{3-cosx}{9-cos^2x} dx\] Not something nice to me :(

OpenStudy (freckles):

wait one sec

OpenStudy (callisto):

Actually, this is the question I wanted to ask. But... start with an easy one first is pretty good.

OpenStudy (anonymous):